1. -4(2x + 1) = -5 - 4x
i.e, -8x-4=-5-4x
i.e, 1=4x
i.e, x=1/4
2. 4(-6x + 2) - 6x = 68
i.e, -24x+8-6x=68
i.e, -30x = 60
i.e, x = -2
3. 3(y - 4) + 6(y + 1) = 57
i.e, 3y-12+6y+6 = 57
i.e, 9y-6 = 57
i.e, y = 7
4. 10 = -2(x - 2) + 6(x - 1)
i.e, 10 = -2x+4+6x-6
i.e, 10 = 4x-2
i.e, x = 3
Please HELP

Answer by Paddock(1)

1. -4(2x + 1) = -5 - 4x
i.e, -8x-4=-5-4x
i.e, 1=4x
i.e, x=1/4
2. 4(-6x + 2) - 6x = 68
i.e, -24x+8-6x=68
i.e, -30x = 60
i.e, x = -2
3. 3(y - 4) + 6(y + 1) = 57
i.e, 3y-12+6y+6 = 57
i.e, 9y-6 = 57
i.e, y = 7
4. 10 = -2(x - 2) + 6(x - 1)
i.e, 10 = -2x+4+6x-6
i.e, 10 = 4x-2
i.e, x = 3